# Designing Speech Technologies for Australian Aboriginal English: Opportunities, Risks and Participation

### v2

@@ 4.4. Evaluation considerations
Doğruöz and Sitaram (2022) have observed that NLP metrics are usually designed with high-resource scenarios in mind. In the case of ASR, the most common evaluation metric is Word Error Rate (WER), which does not actually compare words but rather spelled forms (Karita et al., 2023). It does so in binary fashion: small differences between predictions and reference transcriptions (e.g., ‘naïve’ vs ‘naive’) are penalised just as much as large ones. As such, WER is less useful as a measure of system utility in various contexts, including with language varieties without standardised written forms (Ali et al., 2015; Nigmatulina et al., 2020), and varieties with multiple standard written forms (Kummervold et al., 2024; Karita et al., 2023). In the case of Aboriginal English, WER is not robust to culturally acceptable variation in borrowings from ancestral languages (e.g., ‘Gadigal’ vs ‘Cadigal’ [^4] The Country and people of the Sydney region.) or indigenised spelling variants of English words (e.g., ‘Blak’ vs ‘Black’). [^5] Character Error Rate (CER), another common metric, does not adequately resolve the issue. CER can avoid the overpenalisation of variants accepted by different groups in the community, e.g., only penalising lightly a prediction of ‘Cadigal’ when the reference transcription is ‘Gadigal’ (both refer to the same First Nations Country). However CER is too lenient a metric in cases where the ASR system confuses the spoken word for another word which has a very different meaning or referent, e.g., predicting ‘Kukatj’ when then reference transcription is ‘Kukatja’ (the names of two unrelated Aboriginal language varieties with ancestral lands thousands of kilometres apart). CER would judge such an error as equally bad as the ‘Cadigal’ vs ‘Gadigal’ case.
@@
With a lack of standardised orthography for Aboriginal English, it is important to engage with communities, to centre their perspectives, rather than relying solely on automated metrics such as WER (Cunningham et al., 2024; Markl et al., 2023; Bird, 2020b). As such, the design of the evaluation stage includes interviewing Aboriginal community members, to understand their perspectives on how important fidelity to references is compared to other considerations. An Aboriginal English-specific mapping of commonly accepted spelling variations will be incorporated into WER calculations and enable quantitative measures which provide increased validity with community practices (Ali et al., 2017; Nigmatulina et al., 2020).
4.5. Reflections
We now respond to the questions raised at the start of §3, before providing more specific reflections.
Context.
We acknowledge the colonial context in which this project takes place. Aboriginal and Torres Strait Islander people, the traditional owners of Australia, represent 3.8% of the national population (Australian Government et al., 2020), and are minoritised by the non-Indigenous mainstream population, with significant gaps on economic, educational and health indicators, and life expectancy (Dockery and Milsom, 2007; Commonwealth of Australia, 2024; Australian Institute of Health and Welfare, 2023a, b). Most ancestral languages are extinct or endangered, and Aboriginal English is spoken by around 80% of Aboriginal and Torres Strait Islander people. This project focuses on both Aboriginal English varieties spoken in southwest Western Australia, where Rodríguez Louro and Collard live and work, and varieties of Aboriginal English spoken in the Pilbara (northern Western Australia), Queensland and Alice Springs (Northern Territory) (cf. Risk 2).
Relationships and participation.
The project is a partnership between Google and University of Western Australia. Google is a technology company, many of whose products incorporate ASR technologies. University of Western Australia is an Australian University whose Linguistics department has both expertise in Aboriginal English and existing relationships with Aboriginal communities, including the Nyungar community. Indigenous associates participate in the project, not just as mere providers of voice data, but as partners in the design of elicitation materials and transcription guidelines, management of the voice data collection, and validation and annotation of the collected materials (Opportunity 7). An Indigenous Advisory Committee was also established to provide guidance throughout the project. Investing in these relationships with partners who have earned community trust helps to mitigate risks to relationships with community (cf. Risk 5).
Control.
The project is funded by Google, who set the technological goal of the project: improving ASR for Aboriginal and Torres Strait people in Australia.
Appropriateness.
Prior work by Hutchinson and Cooper has identified that improving ASR for Aboriginal English would bring practical benefits to Indigenous people in Australia (Cooper et al., 2024), by supporting spoken language in many applications (cf. Risk 1).
@@ 4.5. Reflections
We recognise that the legacy of unethical and settler-colonial research has contributed to distrust of research in many Indigenous communities (Smith, 2021). This, along with a distrust of technology companies, can jeopardise even well intentioned projects. Google found that seeking out partnership with (socio)linguists at University of Western Australia with existing relationships, cultural competencies, and sociolinguistic expertise, was critical, and preferable to attempting to build them from scratch (Doğruöz and Sitaram, 2022; Cooper et al., 2024).
@@
Project members took time to reflect on power dynamics, both within the project and in its context (Noe and Kirshenbaum, 2024; Blodgett et al., 2020) (Opportunity 5). We shared and discussed critiques of extractive data practices. We discussed social injustices, including barriers to opportunity for Indigenous people in Australia, and stigmatisation of Aboriginal English. We strengthened accountability through explicit incorporation of ICIP considerations in our research partnership agreement (§7) (cf. Risk 8).
This project aims to go beyond the baseline of ethical language research, and pursue two higher aspirations (Cameron et al., 1993). 1) Advocacy: we treat Aboriginal English as not merely a topic of research but rather aim to support the speaker community (cf. Schwartz, 2022, sec 3.2). This led us to focus on improving ASR, with tangible benefits. 2) Empowerment: we work with the community, using processes that benefit the community independently of project outputs (cf. Cooper et al., 2024; Cunningham et al., 2024).
@@
It was important from the outset to be clear about which Indigenous people and organisations will participate in the project, and how their participation will be facilitated (Schwartz, 2022; Zhang et al., 2022; Lent et al., 2022; Bird, 2020a; Flavelle and Lachler, 2023; Cooper et al., 2024) (cf. Risk 6). In doing so, it was important to be cognisant of the barriers to technology and research careers facing Indigenous participants, to recognise qualifications and authority beyond that of Western institutions, and to recognise contributors as co-authors of research artefacts (Woods, 2023, p. 219; Meakins et al., 2018, pp. 28–29) (cf. Risk 9). While Indigenous participation is critical to the project, we acknowledge that this does not guarantee that the project will be equitable (Noe and Kirshenbaum, 2024; Sloane et al., 2022). We also recognise that the value given to participation and relationships is asymmetric between various project participants. We were also mindful not to force control onto community partners, who are often under-resourced and overburdened (Cameron et al., 1993; Corbett et al., 2023). Establishing the Indigenous Advisory Committee was an important step in helping to negotiate these tensions. We also incorporated aspects of community knowledge and practices into the design process (Winschiers-Theophilus and Bidwell, 2013) (cf. Risk 6), including using an Indigenous-owned consultancy to design culturally appropriate prompts (cf. Risk 7).
4.6. Summary of mitigation of risks and enabling factors of opportunities
@@ 4.7. Limitations
The technological goals of this project were set by Google, and this precluded Indigenous people exercising full control over the project goals. We also acknowledge that Aboriginal and Torres Strait Islander people are less likely to be in full time research positions at companies or universities in Australia, limiting opportunities to be partner investigators on this research.
We acknowledge the plurality of Indigenous cultures and Countries in Australia, and that Indigenous co-investigators, designers, managers and transcribers cannot represent the full diversity of those cultures and Countries. Our decision to include different regions is important not only because of lexico-grammatical differences, but also because Aboriginal English has a key emblematic function as an encoder of different ways of being and doing. We believe that the various ancestries represented wherever Aboriginal English is used are as important as understanding Aboriginal English itself.
Our corpus for developing the initial version of the transcription guidelines was limited to publicly shared YouTube videos. We do not claim that it is representative of spoken Aboriginal English in other domains, such as personal messages, nor of written forms of Aboriginal English.
@@ 5. Discussion: Further Technology Opportunities for Australian Aboriginal English
We now consider technologies supporting Aboriginal English speakers beyond ASR. Given that languages go by multiple names, the NLP community typically uses language codes as identifiers in multilingual systems, datasets, and data catalogues (Kreutzer et al., 2022). This discrete mapping of languages to codes is central to NLP’s conceptual model of what languages are (Smart et al., 2024). Aboriginal English currently lacks a standardised identifier, and this would be useful for future work on this variety. However, given the power of classifying (Bowker and Star, 2000)—and the colonial associations of the origins of the ISO 639-3 identifiers (Dobrin and Good, 2009), which also underpin codes such as the BCP 47—we echo calls for involving community members and linguists, and avoiding offensive exonyms, when designing codes for Aboriginal English (Morey et al., 2013).
Given the primary orality of Aboriginal English, and its distinct grammatical and discourse-pragmatic features relevant to information seeking (§2.1), there are open questions around whether spoken language assistants could better support speakers. Robust support for tag-questions—rather than just questions with subject-verb inversion—might provide more natural support for information requests. While speakers of Aboriginal English benefit from hearing others talk in the same way (Angelo, 2021, p. 42), the question of whether digital assistants should try to ‘sound Indigenous,’ must be approached with care, to avoid potentially reproducing interactions of racialised subservience to non-Indigenous consumers (Owens, 2023).
Before developing writing tools tailored for Aboriginal English, more understanding is required of how writing practices diverge from standardised Australian English. Social media and other domains with informal registers are more likely to have written Aboriginal English, and it may be that modeling orthographic variation is a useful approach for text technologies (cf.  Lin et al., 2024).
Evidence suggests that designing health campaigns in Aboriginal English has community benefits (Collard and Rodríguez Louro, 2025). This suggests potential use cases for machine translation, particularly automated translation and rewriting of health and other social services materials in Aboriginal English.
Finally, given the prevalence of code switching between Aboriginal English and Standardised Australian English, technologies which support the former would benefit from also supporting the latter. Technologies which aim to support English in Australia should explicitly document whether they support Aboriginal English (Bender, 2019).
@@ 6. Conclusions
This paper has explored three questions. First, can speech technology projects support speakers of Australian Aboriginal English by affording specific opportunities? Second, what are the system deployment and procedural risks of such endeavours? Third, how can we mitigate risks, and facilitate opportunities, by integrating participatory practices which appropriately respect the context and are culturally appropriate for the community? We explored these questions through a detailed study of the design considerations of a real-world project that aims to improve ASR for Aboriginal English. The case study of ASR also brings greater attention to a more general trend: the lack of technological support for Indigenous contact language varieties, and the ways that this perpetuates neo-colonialism through a continued failure to recognise and support Indigenous socio-cultural identities and ways of being. Contact language varieties are critical to contemporary Indigenous sociolinguistic identities, and to civil and economic participation, including in places such as Australia where settler-colonialism has led to the extinction of most ancestral languages. We discussed a number of ways in which speech technologies can support speakers of Aboriginal English, as well as the risks of harms caused by system errors. This paper contributes to our understanding of ethical and responsible development of technologies for minoritised language communities, as well as to our understanding of practices for building culturally sensitive technologies while incorporating participatory methods. Although the project focuses on a specific technology, community and context—ASR for speakers of an indigenised variety of English in what we now know as Australia—we believe that some of the considerations and methodologies may be applicable to other technology projects.
@@ 7. Ethical considerations
Before funding was secured for the project, Aboriginal scholars (with expertise in education, applied linguistics, and language revitalisation) were approached and asked to participate as members of a committee advising researchers on the project. Because the research involves Indigenous participants, letters of support from these advisors and from aiatsis (Australian Institute of Aboriginal and Torres Strait Islander Studies) were included in a submission to University of Western Australia’s IRB. Both University of Western Australia’s IRB and Google’s privacy and legal teams reviewed the consent forms that would be used with Aboriginal people creating voice recordings.
In addition, we sought and obtained the support and expertise of the Language Data Commons of Australia (LDaCA) [^6] https://www.ldaca.edu.au/, as the research involves creating a new dataset of Indigenous language materials. LDaCA advised on culturally appropriate data governance and access frameworks (Foley et al., 2024).
The collection of Aboriginal English voice recordings for the purposes of this project raises ICIP issues. As a team, the academic authors at University of Western Australia negotiated with Google the inclusion of an ICIP clause in the research agreement signed in November 2023. This clause reads as follows:
@@
The collection of Aboriginal English for use by Google must be approached with extreme care. From the start, Rodríguez Louro and Collard clarified with their partners that they would not use unstructured yarning sessions like they had done for RA/4/20/4977 ‘Aboriginal English in the global city: Minorities and language change’ (see, e.g., (rodriguez2021working)). With no pre-determined questions and clear Indigenous leadership through the research process, that project collected language as used in everyday settings, including sensitive stories about dispossession, trauma and death by suicide. These materials are firmly closed access with no risk of the audio files and associated transcriptions entering the public domain. However, unstructured yarning sessions tend to give rise to highly culturally sensitive material, which should be protected from use by third parties, especially large technology companies like Google. For this reason, in consultation with Hutchinson and the Advisory Committee, Rodríguez Louro and Collard recommended for this project the design of conversational prompts and story boards (see Appendix A) which seek to elicit interaction on a variety of general topics while steering clear of culturally sensitive material.
Once collected, all data will be screened by a team of Indigenous RAs, led by Collard, and consisting of Narrier and Jones, who will vet the audio files and transcriptions for cultural sensitivity. Any culturally unsafe material (e.g., stories of specific people, places, Indigenous knowledges, trauma, and loss) will be erased from the corpus before sharing the data with others.

### v3

@@ 4.4. Evaluation considerations
Doğruöz and Sitaram (2022) have observed that NLP metrics are usually designed with high-resource scenarios in mind. In the case of ASR, the most common evaluation metric is Word Error Rate (WER), which does not actually compare words but rather spelled forms (Karita et al., 2023). It does so in binary fashion: small differences between predictions and reference transcriptions (e.g., ‘naïve’ vs ‘naive’) are penalised just as much as large ones. As such, WER is less useful as a measure of system utility in various contexts, including with language varieties without standardised written forms (Ali et al., 2015; Nigmatulina et al., 2020), and varieties with multiple standard written forms (Kummervold et al., 2024; Karita et al., 2023). In the case of Australian Aboriginal English, WER is not robust to culturally acceptable variation in borrowings from ancestral languages (e.g., ‘Gadigal’ vs ‘Cadigal’ [^6] The Country and people of the Sydney region.) or indigenised spelling variants of English words (e.g., ‘Blak’ vs ‘Black’). [^7] Character Error Rate (CER), another common metric, does not adequately resolve the issue. CER can avoid the overpenalisation of variants accepted by different groups in the community, e.g., only penalising lightly a prediction of ‘Cadigal’ when the reference transcription is ‘Gadigal’ (both refer to the same First Nations Country). However CER is too lenient a metric in cases where the ASR system confuses the spoken word for another word which has a very different meaning or referent, e.g., predicting ‘Kukatj’ when then reference transcription is ‘Kukatja’ (the names of two unrelated Aboriginal language varieties with ancestral lands thousands of kilometres apart (Australian Institute of Aboriginal and Torres Strait Islander Studies ([n. d.]), AIATSIS)). CER would judge such an error as equally bad as the ‘Cadigal’ vs ‘Gadigal’ case.
@@
With a lack of standardised orthography for Australian Aboriginal English, it is important to engage with communities, to centre their perspectives, rather than relying solely on automated metrics such as WER (Cunningham et al., 2024; Markl et al., 2023; Bird, 2020b). As such, the design of the evaluation stage includes interviewing Aboriginal community members, to understand their perspectives on how important fidelity to references is compared to other considerations. An Australian Aboriginal English-specific mapping of commonly accepted spelling variations will be incorporated into WER calculations and enable quantitative measures which provide increased validity with community practices (Ali et al., 2017; Nigmatulina et al., 2020) (cf. ‘lexeme error rate’ of (Bird, 2021)).
4.5. Designing the project
@@
The project is a partnership between Google and The University of Western Australia (UWA). Google is a technology company, many of whose products incorporate ASR technologies. The University of Western Australia is an Australian University whose Linguistics department has both expertise in Australian Aboriginal English and existing relationships with Aboriginal communities, including the Nyungar community. We recognise that the legacy of unethical and settler-colonial research has contributed to distrust of research in many Indigenous communities (Smith, 2021). This legacy, along with a distrust of technology companies, can jeopardise even well intentioned projects. Google found that seeking out partnership with (socio)linguists at The University of Western Australia with existing relationships, cultural competencies, and sociolinguistic expertise, was critical, and preferable to attempting to build them from scratch (Doğruöz and Sitaram, 2022; Cooper et al., 2024). With regards to control of the project and its outputs (Q3 of §2.5), the project is funded by Google, who set the technological goal of the project: improving ASR for First Nations people in Australia. Google will own any ASR models that they develop using the voice recordings, however UWA will retain ownership of the recordings and serve as stewards of the collection for its appropriate use, by setting up culturally appropriate data governance. We strengthened accountability through explicit incorporation of ICIP considerations in our research partnership agreement (§7) (cf. Risk 8).
Before funding was secured for the project, Aboriginal scholars (with expertise in education, applied linguistics, and language revitalisation) were approached and asked to participate as members of a committee advising researchers on the project. Because the research involves Indigenous participants, letters of support from advisors Professor Clint Bracknell and Mx Sharon Davis, and from aiatsis (Australian Institute of Aboriginal and Torres Strait Islander Studies), were included in a submission to The University of Western Australia’s IRB. Both The University of Western Australia’s IRB and Google’s privacy and legal teams reviewed the consent forms that would be used with Aboriginal people creating voice recordings. In addition, we sought and obtained the support and expertise of the Language Data Commons of Australia (LDaCA), [^8] https://www.ldaca.edu.au/ as the research involves creating a new dataset of Indigenous language materials. LDaCA advised on culturally appropriate data governance and access frameworks (Foley et al., 2024).
The collection of Australian Aboriginal English voice recordings for the purposes of this project raises Indigenous Cultural and Intellectual Property (ICIP) issues. As a team, the academic authors at The University of Western Australia negotiated with Google the inclusion of an ICIP clause in the research agreement signed in November 2023. This clause reads as follows:
Most ancestral languages are extinct or endangered, and Australian Aboriginal English is spoken by around 80% of First Nations people in Australia. This project focuses on both Australian Aboriginal English varieties spoken in southwest Western Australia, where Rodríguez Louro and Collard live and work, and varieties of Australian Aboriginal English spoken in the Pilbara (northern Western Australia), Queensland and Alice Springs (Northern Territory) (cf. Risk 2).
@@
Project members took time to reflect on power dynamics, both within the project and in its context (Noe and Kirshenbaum, 2024; Blodgett et al., 2020) (Opportunity 5). We shared and discussed critiques of extractive data practices. We discussed social injustices, including barriers to opportunity for Indigenous people in Australia, and stigmatisation of Australian Aboriginal English.
@@
It was important from the outset to be clear about which Indigenous people and organisations would participate in the project, and how their participation would be facilitated (Schwartz, 2022; Zhang et al., 2022; Lent et al., 2022; Bird, 2020a; Flavelle and Lachler, 2023; Cooper et al., 2024) (cf. Risk 6). In doing so, it was important to be cognisant of the barriers to technology and research careers facing Indigenous participants, to recognise qualifications and authority beyond that of Western institutions, and to recognise contributors as co-authors of research artefacts (Woods, 2023, p. 219; Meakins et al., 2018, pp. 28–29) (cf. Risk 9). While Indigenous participation is critical to the project, we acknowledge that this does not guarantee that the project will be equitable (Noe and Kirshenbaum, 2024; Sloane et al., 2022). We also recognise that the value given to participation and relationships is asymmetric between various project participants. We were also mindful not to force control onto community partners, who are often under-resourced and overburdened (Cameron et al., 1993; Corbett et al., 2023). Establishing the Indigenous Advisory Committee (see below) was an important step in helping to negotiate these tensions. We also incorporated aspects of community knowledge and practices into the design process (Winschiers-Theophilus and Bidwell, 2013) (cf. Risk 6), including hiring Goorlil Consulting, a Canberra-based consulting firm owned and run by Indigenous experts, to curate and design culturally appropriate visual prompts for the data collection (cf. Risk 7).
An Indigenous Advisory Committee was established to provide guidance throughout the project. The advisors include Professor Clint Bracknell and Mx Sharon Davis. Bracknell is a respected Nyungar academic with a wealth of expertise in projects examining how Indigenous data should be collected and archived for maximum cultural safety. Davis is a renowned education leader from both Bardi and Kija peoples of the Kimberley in Western Australia. Davis has expertise in, and is a champion of, Australian Aboriginal English. Collaborating with Indigenous partners who have expertise in working with community using cultural safety frameworks helps us mitigate risks involving potential harm to Indigenous people and communities (cf. Risk 5). Consultation with the Advisory Committee takes place quarterly. The Advisory Committee meetings, conducted online due to members living in disparate locales across Australia, involve project updates delivered by Rodríguez Louro, followed by Q&A on project progress and outcomes. The Advisory Committee’s role is to provide critical feedback on project conduct and progress. For example, our first meeting involved discussing culturally-safe ways to collect the project data, while additional meetings discussed how to share the collected data, and how to ensure that the bulk of the research assistants for this project are Indigenous.
Indigenous associates participate in the project, not just as mere providers of voice data, but as partners in the design of elicitation materials and transcription guidelines, management of the voice data collection, and validation and annotation of the collected materials (Opportunity 7). All of the research assistants (RAs) conducting fieldwork for this project are First Nations people. While Collard, our senior research assistant, is also a project leader, three additional RAs were also hired. This team includes three Nyungar fieldworkers: Hope Narrier, Katrina Cox and Lily Hayward. While Narrier was introduced to the leadership team through one of Rodríguez Louro’s PhD students, Cox and Hayward were recommended to the project team by UWA’s School of Indigenous Studies. This recommendation was made after Rodríguez Louro contacted the School’s administration with an offer to accept Expressions of Interest for the positions of research assistants, which Cox and Hayward were successful in obtaining.
The maintenance of relationships with Indigenous stakeholders is critical (§2.5). Rodríguez Louro and Collard are in daily contact via phone messaging, a practice that dates back to 2018 when they began their research together. Rodríguez Louro also maintains regular contact with the Indigenous RAs and quarterly direct contact with the Advisory Committee, although contact on social media is more frequent.
Additionally, outreach on the project to Indigenous communities facilitates greater accountability through public visibility. The project has been reported on by Indigenous media outlets (Dudley-Nicholson, 2025; Indigenous Business News, 2025), and Collard has been interviewed about the project on national radio stations (Collard, 2025). Collard, Rodríguez Louro and Hutchinson will also jointly present on aspects of the project at the PULiiMA 2025 Indigenous Languages and Technology Conference in Darwin, Northern Territory, a forum which expects over 700 Indigenous attendees.
@@ 4.6. Limitations
The technological goals of this project were set by Google, and this precluded Indigenous people exercising full control over the project goals. In other words, this project was not driven by or from the community, limiting the opportunities for self-determination. Although innovating on method, and embracing alternate methodologies, the project is nevertheless situated within the dominant paradigm of NLP that seeks to address technical shortcomings through richer datasets, rather than other paradigms which focus on local human agency (Bird, 2024). We also acknowledge that First Nations people in Australia are less likely to be in full time research positions at companies or universities in Australia, limiting opportunities to be partner investigators on this research.
We acknowledge the plurality of Indigenous cultures and Countries in Australia, and that Indigenous co-investigators, designers, managers and transcribers cannot represent the full diversity of those cultures and Countries. Our decision to include multiple regions is important not only because of lexico-grammatical differences, but also because Australian Aboriginal English has a key emblematic function as an encoder of different ways of being and doing. We believe that the various ancestries represented wherever Australian Aboriginal English is used are as important as understanding Australian Aboriginal English itself.
Our corpus for developing the initial version of the transcription guidelines was limited to publicly shared YouTube videos. We do not claim that it is representative of spoken Australian Aboriginal English in other domains, such as personal messages, nor of written forms of Australian Aboriginal English.
@@ 5. Discussion: Further Technology Opportunities for Australian Aboriginal English
We now consider technologies supporting Australian Aboriginal English speakers beyond ASR. Given that languages go by multiple names, the NLP community typically uses language codes as identifiers in multilingual systems, datasets, and data catalogues (Kreutzer et al., 2022). This discrete mapping of languages to codes is central to NLP’s conceptual model of what languages are (Smart et al., 2024; Bird, 2024). Australian Aboriginal English currently lacks a standardised identifier, and this would be useful for future work on this variety. However, given the power of classifying (Bowker and Star, 2000)—and the colonial associations of the origins of the ISO 639-3 identifiers (Dobrin and Good, 2009), which also underpin codes such as the BCP 47—we echo calls for involving community members and linguists, and avoiding offensive exonyms, when designing codes for Australian Aboriginal English (Morey et al., 2013).
Given the primary orality of Australian Aboriginal English, and its distinct grammatical and discourse-pragmatic features relevant to information seeking (§2.1), there are open questions around whether spoken language assistants could better support speakers. Robust support for tag-questions—rather than just questions with subject-verb inversion—might provide more natural support for information requests. While speakers of Australian Aboriginal English benefit from hearing others talk in the same way (Angelo, 2021, p. 42), the question of whether digital assistants should try to ‘sound Indigenous,’ must be approached with care, to avoid potentially reproducing interactions of racialised subservience to non-Indigenous consumers (Owens, 2023).
Before developing writing tools tailored for Australian Aboriginal English, more understanding is required of how writing practices diverge from standardised Australian English. Social media and other domains with informal registers are more likely to have written Australian Aboriginal English, and it may be that modeling orthographic variation is a useful approach for text technologies (cf.  Lin et al., 2024).
Evidence suggests that designing health campaigns in Australian Aboriginal English has community benefits (Collard and Rodríguez Louro, 2025). This suggests potential use cases for machine translation, particularly automated translation and rewriting of health and other social services materials in Australian Aboriginal English.
Finally, given the prevalence of switching between Australian Aboriginal English and Standardised Australian English, technologies which support the former would benefit from also supporting the latter. Technologies which aim to support English in Australia should explicitly document whether they support Australian Aboriginal English (Bender, 2019).
@@ 6. Conclusions
This paper has explored three questions. First, can speech technology projects support speakers of Australian Aboriginal English by affording specific opportunities? Second, what are the system deployment and procedural risks of such endeavours? Third, how can we mitigate risks, and facilitate opportunities, by integrating participatory practices which appropriately respect the context and are culturally appropriate for the community? We explored these questions through a detailed study of the design considerations of a real-world project that aims to improve ASR for Australian Aboriginal English. The case study of ASR also brings greater attention to a more general trend: the lack of technological support for Indigenous contact language varieties, and the ways that this perpetuates neo-colonialism through a continued failure to recognise and support Indigenous socio-cultural identities and ways of being. Contact language varieties are critical to contemporary Indigenous sociolinguistic identities, and to civil and economic participation, including in places such as Australia where settler-colonialism has led to the extinction of most ancestral languages. We discussed a number of ways in which speech technologies can support speakers of Australian Aboriginal English, as well as the risks of harms caused by system errors. This paper contributes to our understanding of ethical and responsible development of technologies for minoritised language communities, as well as to our understanding of practices for building culturally sensitive technologies while incorporating participatory methods. Although the project focuses on a specific technology, community and context—ASR for speakers of an indigenised variety of English in what we now know as Australia—we believe that some of the considerations and methodologies may be applicable to other technology projects.
@@ 7. Ethical considerations
The collection of Australian Aboriginal English for use by Google must be approached with extreme care. From the start, Rodríguez Louro and Collard clarified with their partners that they would not use unstructured yarning sessions like they had done for RA/4/20/4977 ‘Aboriginal English in the global city: Minorities and language change’ (see, e.g., (Rodríguez Louro and Collard, 2021b)). With no pre-determined questions and clear Indigenous leadership through the research process, that project collected language as used in everyday settings, including sensitive stories about dispossession, trauma and suicide. These materials are firmly closed access with no risk of the audio files and associated transcriptions entering the public domain. However, unstructured yarning sessions tend to give rise to highly culturally sensitive material, which should be protected from use by third parties, especially large technology companies like Google. For this reason, in consultation with Hutchinson and the Advisory Committee, Rodríguez Louro and Collard recommended for this project the design of conversational prompts and story boards (see Appendix A) which seek to elicit interaction on a variety of general topics while steering clear of culturally sensitive material.
Only publicly shared YouTube videos were included in the corpus for developing transcription guidelines, and the data was not copied, modified or re-published, however the video creators were not contacted in order to understand their sentiments towards this research.
8. Positionality
Collard, Hutchinson, and Cooper are born in Australia, and Rodríguez Louro is an Australian born and raised in Argentina. Together, we have ancestral connections to Australian Indigenous Countries, Europe and South America, and we have trained and/or worked in linguistics, education, law, natural language processing, software engineering, and cybernetics. Collard is an L1 speaker of Australian Aboriginal English, and Collard and Rodríguez Louro have been active in advocating for greater appreciation of Australian Aboriginal English in civil society.
Acknowledgements.
The University of Western Australia, where the second and third authors are based, is situated on unceded Nyungar Country; Google’s Sydney office, where the first author is based, is situated on unceded Gadigal Country; and The Australian National University’s Acton campus, where the last author is based, is situated on unceded Ngunnawal and Ngambri Country. All authors acknowledge that the Nyungar, Gadigal, Ngunnawal and Ngambri people respectively remain the spiritual and cultural custodians of their Country, on which they continue to practise their values, languages, beliefs, and knowledge. We would like to thank the Advisory Committee members Prof. Clint Bracknell and Mx Sharon Davis for their guidance; research assistants Hope Narrier, Katrina Cox and Lily Hayward for their data collection with members of their community; Emily Gilchrist for her help with project administration and data processing; Nefeli Perdikouli, Mitch Browne, Grace Shepherd, Madeleine Clews and Lucia Fraiese for their help identifying Aboriginal English in YouTube videos; and Vera Axelrod, Daan van Esch, Prof. Michael Haugh, Robert McLellan and Scott Riddle for their project support and advice. We would also like to thank Steven Bird and the anonymous reviewers of FAccT and other venues for their questions and suggestions.